### v4

# CLUE-Mark: Watermarking Diffusion Models using CLWE

Abstract.
The proliferation of AI-generated images has led to a new problem known as “AI Slop”: low-effort, mass-produced AI-generated images that are distributed as-is but are difficult to distinguish from authentic images. Watermarking such content from popular diffusion models provides a promising solution: if social media platforms and end users can easily identify images generated by the most common AI providers, then a very large proportion of the AI Slop problem can be detected by end users. However, prior watermarking techniques are heuristic and lack formal guarantees of undetectability. Techniques that were claimed to not affect image quality were later shown to in fact degrade quality, hampering adoption as providers and users are averse to any reduction in output quality.
In this work, we introduce CLUE-Mark, a provably undetectable watermaking scheme for diffusion models. CLUE-Mark requires no changes to the model being used, is computationally efficient, and guaranteed to have no impact on model output quality. Our approach leverages the Continuous Learning With Errors (CLWE) problem — a cryptographically hard lattice problem — to embed hidden messages in the latent noise vectors used by diffusion models. We reduce the security of CLWE against an adversary given limited samples to the security of CLUE-Mark. Assuming that CLWE is secure in the limited sample regime, image watermarks generated by CLUE-Mark are undetectable by any efficient adversary that does not have the secret key, making them imperceptible to human observers and undetectable by ad-hoc heuristics as well. CLUE-Mark focuses primarily on undetectability while maintaining sufficient robustness for common non-adversarial transformations such as JPEG compression. Empirical evaluations on state-of-the-art diffusion models confirm that CLUE-Mark achieves high message recovery, preserves image quality, and is robust to minor perturbations such as JPEG compression and brightness adjustments. A proof-of-concept implementation of CLUE-Mark and source code to reproduce our evaluations is available at: https://github.com/kisp-nus/cluemark
1. Introduction
The internet has recently witnessed an unprecedented surge of AI-generated content, an effect that was first popularly referred to as “AI Slop” in mid-2024 (Hern and Milmo, [n. d.]; Hoffman, [n. d.]). This phenomenon has now entered popular discourse with coverage from diverse media outlets including The Guardian (Hern and Milmo, [n. d.]; Malik, 2025; Mahdawi, [n. d.]), The New York Times (Hoffman, [n. d.]), New York Magazine Intelligencer (Read, [n. d.]), and John Oliver’s Last Week Tonight (Oliver, 2025). As pointed out by many commentators, the problem with such content is not the generation of content itself, but that for many users it is difficult to distinguish authentic and synthetic content, leading to problems of misinformation and user trust.
Such machine-generated content is created in large volumes and shared through social media with little to no modification. Many users feel either frustrated by the inability to distinguish it from authetic content or overwhelmed by the sheer volume of it, or both, with many blaming social media platforms for doing too little to identify AI generated content (Hern and Milmo, [n. d.]; Malik, 2025). Thus, the problem of identifying AI-generated images is both timely and important.
One proposed method to identifying AI-generated images is watermarking, in which the image is “marked” in such a way that it can be distinguished from an unmarked image. In this manner, should publicly available AI image generation tools embed watermarks in their output, as has been proposed by several legal projects (ai_, 2023; of Public Affairs, 2024; Diane and Krystal, 2024; Madiega, 2023), it would be possible to identify AI-generated content automatically either by social media platforms or by end-user software. As the name and physical analogy suggests, watermarking is typically designed to be robust to benign transformations, but is also easily detectable by the user, producing at least some visible artifacts or other degradation in image quality. This presents a major barrier to adoption among both providers and users of AI image generation tools, as they are (understandably) reluctant to degrade the quality of the images generated by their models after spending considerable effort to train them.
Prior approaches to watermarking AI-generated images were ad-hoc. For example, Tree Ring (Wen et al., 2023) was conjectured to be undetectable and robust, but no proof as such was given. While an important step forward, Tree Ring watermarks can be both detected and removed by standard steganalysis (Yang et al., 2024a), i.e. averaging the difference between watermarked and non-watermarked images. Similarly, Gaussian Shading (Yang et al., 2024b) was proved to be undetectable, but only under the assumption that the nonce used to generate the watermark is not reused; otherwise the watermark is again detectable by steganalysis. The one-time (per image) nonce assumption is hard to satisfy practically. This is more than a security concern; it shows that the watermark affects the output of the model such that the output images are both quantitatively and qualitatively different from the original (a finding that we verify in Section 5.5).
Thus for both security and quality reasons, it is important that the watermark be provably undetectable, by which we mean that a watermarked image must be indistinguishable from an unwatermarked image111 Embedding a signal in an image such that it cannot be detected but is not robust to perturbations is technically considered steganography rather than watermarking, but but the latter term remains popular.. A rigorous reduction from a cryptographically hard problem is needed, otherwise it may turn out that a watermark that was conjectured to be undetectable is in fact detectable by some other means, or may affect downstream applications. By proving undetectability, we ensure that the watermark does not compromise the quality of the generated images and remains stealthy to detectors. In an ideal world, such an undetectable watermark (properly: steganography) would also be robust to benign transformations. Unfortunately, it has been proved that a perfect watermark cannot exist (see Appendix A and (Zhang et al., 2024)), and thus we must choose between robustness and undetectability.
In this work we present CLUE-Mark, an undetectable watermark (i.e. steganography) scheme for diffusion models to address the problem of AI Slop. Given the challenges listed above and the impossibility of perfect watermarking, we focus on using cryptographic techniques to inject hidden information into diffusion model generated images such that any efficient algorithm that can distinguish marked and unmarked images can be used to break fundamental cryptographic assumptions — something assumed to be impossible — while verifiers that have access to the secret key can recover the hidden information with high probability, despite minor purturbations such as JPEG compression. This allows labelling of AI-generated content at scale, precise the problem posed by AI Slop. Even if sophisticated users employing more effort or advanced techniques can remove the hidden watermark, this prevents low-effort users from flooding social media networks with unmarked AI-generated content, drastically improving the scale of the problem. By proving undetectability, we ensure that our method does not compromise the quality of the generated images.
Undetectability from CLWE
Our approach modifies the noise input to a diffusion model using a distribution based on the Continuous Learning With Errors (CLWE) problem (Bruna et al., 2021). Intuitively, this distribution is gaussian (i.e. normal) in all directions except for a secret direction in which it has a signature pattern. For our CLUE-Mark construction, we can prove that any process that can reliably detect this hidden pattern can be used to distinguish a CLWE distribution from a normal distribution. Assuming that this problem is cryptographically hard, we can guarantee that our approach is undetectable by any efficient algorithm without the secret key, ensuring perfect image quality and compatibility with all content creation tools. Because our technique only modifies the input noise, it does not require any model modifications or retraining.
Empirical Results
While CLUE-Mark aims for undetectability, we must also demonstrate its practical utility. Can our steganographic signal be reliably recovered after image generation and common non-adversarial transformations like JPEG compression? To answer this, we evaluate CLUE-Mark on commonly used diffusion models and datasets to verify undetectability, recoverability, image quality, and robustness to non-adversarial transformations. We show that CLUE-Mark is undetectable by standard means, produces images with identical quality to those without steganographic content, and the hidden message can be reliably recovered after JPEG compression and minor brightness adjustments. A proof-of-concept implementation of CLUE-Mark and source code to reproduce our evaluations is available at: https://github.com/kisp-nus/cluemark
Comparison to PRC-based Watermarks
In work contemporary to ours, Gunn et al (Gunn et al., 2025) propose using Pseudo-Random Codes (PRC) to bias the latent input to a diffusion model for watermarking. While PRC offer a promising new primitive, our work uses CLWE—a complementary primitive to encode secret information in the continuous $\mathbb{R}^{d}$ domain directly—which is more natural fit for machine learning systems. PRC-based watermarks use parameters that are not known to be theoretically secure, as stated in their work, and is more suitable for applications where robustness to adversarial perturbation is more important than undetectability. We discuss in more detail how our work differs from PRC watermarks in Section 6.
Contributions.
Our contribution is a novel and practical watermarking scheme, CLUE-Mark, for images generated by modern diffusion AI models. It aims to be undetectable, in the sense that distinguishing marked and unmarked images is as hard as distinguishing the CLWE distribution from a Gaussian. This assumption is stricter than the standard cryptographic hardness of CLWE; it considers the practically-motivated regime where the adversary has limited samples. Along the way we develop a novel hCLWE sampling algorithm (see Section 4.3) that is more practical than the previous rejection-sampling technique proposed in (Bruna et al., 2021), which may be of independent interest. These tools enables detection of low-effort AI-generated content (“AI Slop”) without compromising image quality or compatibility with existing tools and platforms.
2. Motivation and Problem Setup
Figure 1. Sampling images from latent diffusion models
Diffusion models, such as Stable Diffusion (Podell et al., 2023; Rombach et al., 2022), have quickly emerged as the go-to AI models for generating high quality and customizable images. They have been adopted by content creators across diverse fields, from digital art to professional illustration for e-commerce and advertising (see survey (Yang et al., 2023)). The process to generate an image from a diffusion model is called “sampling,” in which the model is given a text embedding and a noise vector (in the latent space of the model, hence this is often called the “latent vector” or “latent noise”), and the model iteratively removes the noise to produce an image described by the text embedding, as illustrated in Figure 1. We refer to (Rombach et al., 2022; Podell et al., 2023) for understanding the details of diffusion models, and describe only the parts that are necessary for our work.
Definition 2.0 (Diffusion Model Sampling).
The process to sample a diffusion model is defined as the function:
$\text{Sample}_{\theta}(\pi\in\mathcal{T}^{*};\mathbf{z}\in\mathbb{R}^{n})\rightarrow\mathbf{x}\in\mathbb{R}^{N}$
where $\theta$ is the model parameters, $\pi$ is the embedding of the text prompt, $\mathbf{z}$ is the noise vector, and $\mathbf{x}$ is the generated image. The noise vector is drawn from a standard normal (Gaussian) distribution, that is $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
Watermarking Diffusion Models.
Watermarking is a proposed measure to detect AI-generated content, which can be used to fight against AI Slop. Watermarking an image refers to implanting a signal within it such that the signal can later be recovered, providing proof of the provenance of the image. In a secret key scheme, only those with the key can verify an image’s origin, which is technically closer to steganography than watermarking, the latter usually prioritizes robustness to perturbations over undetectability. Consider a company that hosts a closed-source model. If the company watermarks the images generated by the model using a secret key, they can later verify whether an image was generated by the model by checking for the watermark using the key. They can also give the key to another entity to do verification, for example a social media platform that wants to verify the provenance of images uploaded by users. The company may also assign a key to each user in order to trace which user was responsible for generating a particular image. Notice that the verifier, whether the model provider, social media platform, or another party, is given the model keys but not any per-image data, such as text prompts, nonces, or noise vectors. Watermarks may be applied to images after generation (“post-process”) or during model execution (“in-process”). Diffusion models offer another option: watermarking the noise vector $\mathbf{z}$ used to generate the image (“pre-process”), as introduced by Tree Rings (Wen et al., 2023). In this work we focus on secret-key pre-process watermarking. See Figure 2 for an illustration of the use cases.
Figure 2. Watermarking use cases and our threat model, with a per-model (and per-creator) keys. Note that the verifier is given model keys, but no image-specific data, such as text prompts, noise vectors, or per-image nonces or keys.
Previous approaches
Despite the potential benefits of watermarking AI generated images, adoption has been limited (Seetharaman and Barnum, 2023). One reason for this is image quality. Large models that generate high-quality images are expensive to train and serve, and both companies and users are sensitive to any degradation in image quality. Further, changes to image quality may also affect the performance of downstream tasks. For example, if the generated images are used to train another model or as inputs to a larger process, any changes in the images may affect the performance of the system as a whole.
Existing watermarking techniques for diffusion models have a noticeable impact on the images generated. Post-processing techniques, e.g. embedding a signal into the output pixel data and even AI-based techniques, have visible quality losses (Wen et al., 2023). As we evaluate in this work, state-of-the-art techniques such as Tree Ring and Gaussian Shading, despite aiming for not being perceptible, produce images with a significant quality differences from their unwatermarked equivalents (see Section 5.4). This is because of ad-hoc design choices and the lack of provable guarantees. When evaluated over a large set of images, it can be seen that the watermarked images are biased (less diverse statistically) than the unwatermarked images. Further, recent attacks have shown that the watermarks can be detected, removed, and forged using standard steganographic techniques (Yang et al., 2024a).
Problem
The question thus becomes: can we watermark diffusion model generated images such that a verifier as in Figure 2 can distinguish between watermarked and unwatermarked images, while proving that the watermark does not affect the quality of the images? We formally define this problem in the next section.
2.1. Undetectable Watermarking
In this work, we aim to solve this problem by designing provably undetectable watermarking. By this we mean that distinguishing between watermarked and unwatermarked images must be computationally at least as hard as solving a cryptographic problem. This ensures that there is a negligible difference between watermarked and unwatermarked images for any downstream application. Even a determined adversary trying just to detect the presence of a watermark faces a challenge at least as hard as breaking state-of-the-art cryptography. This approach embodies a secure by construction principle, in contrast to existing techniques which empirically evaluate against specific attacks but cannot rule out the existence of other effective attacks. In order to construct a provable scheme, we first formally define the watermarking process.
Definition 2.0 (Watermarking).
A watermarking function is defined as the tuple of functions $(\text{Setup},\text{Mark}$, $\text{Extract})$:
$\text{Setup}_{\theta}(1^{\lambda})\rightarrow k\in\mathcal{K}$
$\text{Mark}_{\theta}(\pi\in\mathcal{T}^{*};k\in\mathcal{K})\rightarrow\mathbf{x}\in\mathbb{R}^{N}$
$\text{Extract}_{\theta}(\mathbf{x}\in\mathbb{R}^{N};k\in\mathcal{K})\rightarrow\{0,1\}$
where $\lambda$ is a security parameter, $\theta$ is the model parameters, $\pi$ is the embedding of the text prompt, $k$ is the secret key, and $\mathbf{x}$ is the generated image.
For a watermarking scheme to be useful, it needs to satisfy both completeness and soundness criteria, for which we use Definitions 3 and 4 below. Notice that the definitions are worst-case over all prompts. Our approach currently focuses on embedding a single bit of information, though this can be extended to multiple bits through phase-shifting techniques or using multiple keys.
Definition 2.0 (Watermark Completeness).
A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is $\delta$-complete if for all $\pi\in\mathcal{T}^{*}$:
$\Pr\left[\text{Extract}_{\theta}(\text{Mark}_{\theta}(\pi;k);k)=1|k\leftarrow\text{Setup}_{\theta}(1^{\lambda})\right]\geq 1-\delta$
Where the probability is taken over all random coins used in Setup, Mark, and Extract.
Definition 2.0 (Watermark Soundness).
A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is $\delta$-sound if for all $\pi\in\mathcal{T}^{*}$:
$\Pr[\text{Extract}_{\theta}($ $\text{Sample}_{\theta}(\mathbf{z};\pi);k)=1|$
$k\leftarrow\text{Setup}_{\theta}(1^{\lambda}),\mathbf{z}\leftarrow\mathcal{N}(\mathbf{0},\mathbf{I})]\leq\delta$
Where the probablity is taken over $\mathbf{z}$ and all random coins used in Setup.
Before we can define undetectability, we must define formally what we mean by “distinguishing” and how we measure success. For this we use Definition 5.
Definition 2.0 (Advantage of a Distinguisher).
Let $D^{\mathcal{O}}\rightarrow\{0,1\}$ be a distinguisher with access to oracle $\mathcal{O}$. Define the advantage of $D$ in distinguishing $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ as:
$Adv_{D,\mathcal{O}_{1},\mathcal{O}_{2}}=\left|\Pr\left[D^{\mathcal{O}_{1}}(1^{\lambda})=1\right]-\Pr\left[D^{\mathcal{O}_{2}}(1^{\lambda})=1\right]\right|$
We want to ensure that our watermarking scheme is undetectable, that is, that the output of the watermarking process is indistinguishable from non-watermarked images generated by the model. For both users and providers, this guarantees that watermarked outputs have identitical quality to unmarked outputs, and remains compatible with all existing tools and platforms. We formally define this in Definition 6.
Definition 2.0 (Provable Undetectability).
Let $\mathcal{M}_{\theta}$ be an oracle that on input $\pi\in\mathcal{T}^{*}$ samples $\mathbf{z}\leftarrow\mathcal{N}(\mathbf{0},\mathbf{I})$ and returns $\text{Sample}_{\theta}(\pi;\mathbf{z})$. Let $\mathcal{S}_{\theta,k}$ be an oracle that on input $\pi\in\mathcal{T}^{*}$ returns $\text{Mark}_{\theta}(\pi;k)$ for a fixed key $k\leftarrow\text{Setup}_{\theta}(1^{\lambda})$. A watermarking scheme $(\text{Setup},\text{Mark}$, $\text{Extract})$ is undetectable if for all probabilistic polynomial-time PPT distinguishers $D$, $Adv_{D,\mathcal{M}_{\theta},\mathcal{S}_{\theta,k}}\leq negl(\lambda)$.
While robustness against adversarial manipulations is a secondary concern in our AI Slop prevention use case, we still require the embedded information to survive common non-adversarial transformations that occur in normal usage, such as JPEG compression when images are shared on social media platforms (see Figure 2). Our approach prioritizes absolute undetectability first, while maintaining sufficient robustness for practical usage.
3. The CLUE-Mark Approach
Our approach to steganography is based on the idea of modifying the latent noise vector of a diffusion model (i.e. pre-process steganography). Recall that a diffusion model takes as input a latent vector that is expected to be from the standard gaussian distribution. In our approach, we instead use a latent vector that is drawn from the hCLWE distribution (which will be defined in Section 3.1). This vector is indistinguishable from the standard gaussian distribution to all PPT-computable functions without the secret key, but can be efficiently distinguished when the secret key is known. We call using hCLWE samples to generate an image the signal injection process. To test whether an image contains a signal, a system with access to the secret key uses the signal recovery process. During recovery, we invert the diffusion model to get an estimate of the initial latent vector from the image. We then use the secret key to determine if the latent vector is from the hCLWE distribution, and if so, then it is very likely that the image contains a signal using that key.
In this section we describe the high-level approach of steganography, including the necessary background on CLWE, and the signal injection and recovery processes. We will prove undetectability of the construction in Section 4.
3.1. Background on CLWE
Consider the scatter plot shown in Figure 3. If you were to look at the distribution of the points along either axis, it would look like a standard gaussian distribution. But if you were to plot the points along the line y=xy=-x, you would notice that the points are periodic. That is, that they’re bunched together around 1,1/2,0,1/2,1-1,-1/2,0,1/2,1, and so on. In two dimensions you can see this immediately by looking at the scatter plot, but what if the points are in three dimensions? The points would form a sphere from nearly any angle you look at, except if you found the “secret direction” you would notice that the points form slices and are again periodic [^2] The original CLWE presentation at https://youtu.be/9CB4KcoRB9U?t=60 has an animation that illustrates this effect clearly..
This is the essence of the homogenous Continuous Learning With Errors (hCLWE) problem [^3] Homogenous CLWE is a special case of the more general CLWE problem.. Intuitively, it becomes harder to find the secret direction as you increase the number of dimensions, as the number of directions you have to check grows exponentially. Deciding whether a set of points (from now on let us call them samples) are normally distributed in all directions, or if there is a secret direction in which they are periodic has been proven to be as hard as solving worst-case lattice problems (Bruna et al., 2021).
More formally, consider a unit vector $\mathbf{w}\in\mathbb{R}^{n}$ chosen uniformly at random, and a set of samples $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\in\mathbb{R}^{n}$ that are drawn from a standard gaussian distribution in all directions except for the direction $\mathbf{w}$, in which case the distribution is shown in Figure 4. This produces the well-known “gaussian-pancakes” distribution, which produces periodic slices with separation $\approx 1/\gamma$ and width $\approx\beta/\gamma$ in the secret direction $\mathbf{w}$, and standard gaussian in all others. The hCLWE problem is to determine the secret direction $\mathbf{w}$ given a polynomial number of samples ${\left\{\mathbf{y}_{i}\right\}}_{i\in[m]}$. The decision variant of hCLWE is to distinguish between samples from the hCLWE distribution and from the standard multivariate normal. This leads to Theorem 1, an informal statement of the CLWE hardness theorem; see (Bruna et al., 2021) for the more formal statement, proof, and further details.
Theorem 1 (hCLWE Hardness Informal).
For $\gamma=\Omega(\sqrt{n}),\beta\in(0,1)$ such that $\beta/\gamma$ is polynomially bounded, if there exists an algorithm that can efficiently distinguish the $\text{hCLWE}_{\gamma,\beta}$ distribution from the standard multivariate gaussian distribution, then there exists an efficient algorithm that approximates worst-case lattice problems to within a factor polynomial in $n$.
Such lattice problems are considered cryptographically hard for appropriately chosen parameters $\gamma$ and $\beta$. Assuming that no PPT algorithm can solve them, then we can state that no PPT algorithm can solve the decision hCLWE problem with non-negligible probability. Assumption 2.1 given later in Section 4.1 is the more precise version we rely on. It constraints the number of samples available.
Figure 3. Example scatter plot of $10,000$ samples from the hCLWE distribution with secret direction $(1/\sqrt{2},-1/\sqrt{2})$, $\gamma=2,\beta=0.1$.
Figure 4. The hCLWE PDF with $\gamma=2,\beta=0.1$ in the secret direction rescaled for comparison with a standard gaussian.
3.2. Creating Watermarks from CLWE
Figure 5. CLUE-Mark signal injection process showing how the latent vector is divided into blocks and transformed using the inverse discrete wavelet transform.
There are several technical challenges we overcome in making CLWE work reliably for diffusion models, which are explained next.
Blocking.
In order to use samples from the hCLWE distribution as inputs to a diffusion model, we must first overcome a small disparity: The hCLWE distribution assumes a number of high-dimensional samples, while the noise vector required for the initial diffusion model latents is a single vector. The solution is to divide the latent vector into a number of non-overlapping blocks, each of which is treated as a sample from the hCLWE distribution. The blocks must each be large enough to ensure undetectability, but small enough to ensure enough samples to recover the signal even if some blocks are lost or corrupted by noise. As we will see in Section 5.1, for a latent vector of dimensions $4\times 64\times 64$ we choose a block size of $2\times 4\times 4$ for our evaluations, but for the moment we leave this as an implementation parameter.
Frequency Domain.
Using hCLWE samples directly does not give robust recovery of the secret embedded. In order to improve robustness to perturbations and noise during recovery, we use a standard technique in image watermarking (see (Al-Haj, 2007)): we apply the hCLWE samples in the frequency domain of the latents, i.e., we take the inverse discrete wavelet transform (IDWT) of the hCLWE samples to arrive at the latent vector used to generate the image. The process for image generation is shown in Figure 5. The signal recovery process reverses this process, but requires additional steps in order to extract the signal in the presence of noise and perturbations.
Recovery.
We now have an approach that allows us to embed an hCLWE distribution in the latent space of a diffusion model, which we then use to produce an image. The question becomes: how do we determine if an image contains a signal? From recent work (Hong et al., 2024; Song et al., 2020), we can invert the diffusion model to get an estimate of the original latent vector (assuming the image is not perturbed)—even without the original prompt or embedding vector used to create the image, albeit with some noise. Dividing the latent vector into blocks is then trivial, but the resulting blocks are subject to noise and perturbation. Given these noisy blocks and the secret direction of the hCLWE distribution, how do we determine if the secret signal is present?
Figure 6. Signal recovery process showing how the latent vector is extracted from the image, divided into blocks, and analyzed using the secret direction.
First, we can measure the error from the ideal hCLWE distribution by computing the inner product with the secret direction to get a set of $z_{i}$ values:
$z_{i}=\gamma\langle\mathbf{y}_{i},\mathbf{w}\rangle\bmod 1$
In the no-signal case, we expect that for well-chosen parameters the $z_{i}$ values will be uniformly distributed in $[0,1)$. In the signal case, we expect that the $z_{i}$ values will be concentrated near zero, with the strength of the concentration depending on $\beta$, noise, and perturbations. This produces the situation shown in Figure 7, which shows the result of a statistical simulation comparing $z_{i}$ values calculated from samples drawn from a standard normal, hCLWE, and hCLWE with noise from a normal distribution of width $0.2$.
Figure 7. Rose diagram of $z_{i}$ values from statistical simulation of $10,000$ samples showing the distribution of errors for different cases, where a $z$ score of $1.0$ is mapped to $360\degree$.
We now need to decide between two hypotheseses: the null hypothesis is that the samples are from a standard normal distribution, and thus the $z_{i}$ values are uniformly distributed in $[0,1)$, while the alternative hypothesis is that the samples are from the hCLWE distribution, and thus the $z_{i}$ values have some a concentration around some (unimodal) peak. Fortunately, the Rayleigh test (Ruxton, 2017) is a well-known test in circular statistics that provides the optimal ability to distinguish between these two cases, and provides a $p$-value that indicates the probability that the null hypothesis is true given the observed data. Thus, if the $p$-value is below a certain threshold, we can reject the null hypothesis and conclude that the samples are from the hCLWE distribution, and thus the image contains a signal. For example, in the statistical simulation shown in Figure 7, if we restrict to the first 256 samples (as we would expect in our use case), the Rayleigh test gives a $p$-value of $36.5\%$ for the standard normal samples, $1.2\times 10^{-94}$ for the hCLWE samples, and $5.2\times 10^{-69}$ for the hCLWE samples with noise.
The resulting recovery process is shown in Figure 6. In terms of parameters, we set $n$ (the number of dimensions per CLWE sample), $\gamma$ and $\beta$ (CLWE parameters) for security of the reductions in Section 5.1, and $m$ is the size/dimensions of the latent vector divided by $n$.
4. Security Proof
In this section, we provide the formal definitions and proofs of completeness, soundness, and undetectability of CLUE-Mark.
4.1. Formal Definitions
Define:
(1) $\rho_{s}(\mathbf{x})=\exp\left(-\pi\|\mathbf{x}/s\|^{2}\right)$
Where $\mathbf{x}\in\mathbb{R}^{n}$. Note that $\rho_{s}(\mathbf{x})/s^{n}$ is the PDF of the gaussian distribution with covariance $s^{2}/2\pi\cdot I_{n}$.
Definition 4.0 (hCLWE Distribution).
Let $\mathbf{w}\in\mathbb{R}^{n}$ be a unit vector, and $\beta,\gamma>0$ be parameters. Define the homogenous continuous learning with errors (hCLWE) distribution $H_{\mathbf{w},\beta,\gamma}$ over $\mathbb{R}^{n}$ to have density at point $\mathbf{y}$ proportional to:
$\rho(\mathbf{y})\cdot\sum_{k\in\mathbb{Z}}\rho_{\beta}(k-\gamma\langle\mathbf{w},\mathbf{y}\rangle)$
Definition 4.0 (Decision hCLWE Problem).
Let $\mathbf{w}\in\mathbb{R}^{n}$ be a unit vector drawn uniformly at random, and $\beta,\gamma>0$ be parameters. Let $\mathcal{N}$ be an oracle that responds to queries with samples from $\mathcal{N}(0,I_{n})$, and let $\mathcal{H}_{\mathbf{w},\beta,\gamma}$ be an oracle that responds with samples from $H_{\mathbf{w},\beta,\gamma}$. The decision hCLWE problem is to distinguish $\mathcal{N}$ and $\mathcal{H}_{\mathbf{w},\beta,\gamma}$. That is for distinguisher $D$, we say that $D$ solves the decision hCLWE problem with advantage $Adv_{D,\mathcal{N},\mathcal{H}_{\mathbf{w},\beta,\gamma}}$.
Assumption 2.1 (hCLWE).
No PPT distinguisher $D$ can solve the decision hCLWE problem with non-negligible advantage when given a set of $m$ samples, where $m$ is at most polynomial in $n$.
The above assumption is stricter than the standard CLWE hardness, as it considers an adversary with a limited number of samples.
4.2. Undetectability
Theorem 3 gives our main theoretical result: detecting the watermark is as hard as solving the decision hCLWE problem. Before we dive into the technical details of the proof, we first provide an intuitive explanation. Our proof is by reduction, that is if there exists a distinguisher that can reliably distinguish watermarked and unwatermarked images, then we can use that distinguisher to solve the decision hCLWE problem. Recall that in the hCLWE problem the distinguisher is given an oracle that for all queries returns samples from either the hCLWE distribution or the standard normal distribution. In the reduction, whenever the watermark distinguisher requests an image, we construct the latent vector from samples from the given oracle, which we give to the latent diffusion model to generate an image, before passing the image back to the watermark distinguisher. If the oracle gives us standard normal samples, then this corresponds to the unwatermarked case, while if the oracle gives us hCLWE samples, then this corresponds to the watermarked case. Other than the samples used to build the initial latent vector, the process is identical and thus the difference in probability between the two cases can only come from the samples. Note that if the latent vector has dimension $mn$, then each image generated requires $m$ samples from the oracle. If the watermark distinguisher has a non-negligible advantage, then so does our hCLWE distinguisher.
Theorem 3.
Assuming the hCLWE assumption holds for $n$ dimensions with parameters $\gamma,\beta$, the CLUE-Mark scheme described in Section 3 with latent vectors of dimension mnmn such that $m$ is at most polynomial in $n$ is undetectable by Definition 6.
Proof.
Consider a PPT distinguisher $D^{\mathcal{O}}$ that has non-negligible advantage $\epsilon$ in distinguishing between watermarked and non-watermarked images, that is $Adv_{D,\mathcal{M}_{\theta},\mathcal{W}_{\theta,k}}\geq\epsilon$.
We now construct $D^{\prime}$ that solves the decision hCLWE problem, given access to $\mathcal{O}^{\prime}$, which is either $\mathcal{N}$, an oracle that returns samples from the standard normal, or $\mathcal{H}_{\mathbf{w},\beta,\gamma}$, an oracle that gives samples from the hCLWE distribution [^4] The oracle is consistent for all queries, that is if the oracle is $\mathcal{N}$ then all results are standard normal. The oracle does change distributions between queries.. Distringuisher $D^{\prime}$ works as follows:
  1. (1)
    On receiving a request to $\mathcal{O}$ with query $\pi$, $D^{\prime}$:
    1. (a)
      requests $m$ samples from the oracle, $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\leftarrow\mathcal{O}^{\prime},\mathbf{y}_{i}\in\mathbb{R}^{n}$, where $mn$ is the size of the latent space of the model.
    2. (b)
      Assembles the samples to form the dimensions of the latent space, i.e. $\mathbf{z}^{\prime}\leftarrow(\mathbf{y}_{1},\ldots,\mathbf{y}_{m})\in\mathbb{R}^{mn}$.
    3. (c)
      Applies the inverse DWT to the assembled samples to obtain a latent vector, i.e. $\mathbf{z}\leftarrow\text{IDWT}(\mathbf{z}^{\prime})$.
    4. (d)
      Uses the result as the latent vector to sample an image from the model, $\mathbf{x}\leftarrow\text{Sample}_{\theta}(\pi,\mathbf{z})$.
    5. (e)
      Forwards the image $\mathbf{x}$ to $D$.
  2. (2)
    $D^{\prime}$ outputs the same result as $D$.
If $D$ has advantage $\epsilon$ in distinguishing watermarked and unwatermarked images, then $D^{\prime}$ has at least $\epsilon$ advantage in distinguishing hCLWE samples from standard normal samples. Since all of the operations in $D^{\prime}$ are polynomial time, this contradicts the hCLWE assumption. ∎
4.3. Sampling from hCLWE
Sampling from the hCLWE distribution is non-trivial. In (Bruna et al., 2021) it is suggested that the hCLWE distribution be sampled by rejection-sampling from the standard normal distribution. However, this is not practical for our setup, as not only does this require drawing many more samples than needed, it also assumes the base distribution is the standard normal and easy to sample. In our case, the base distribution is the model’s latent space after a DWT transform. Rather than rejection sampling in this space, we propose a much simpler approach: beginning with samples of the appropriate base distribution (e.g. standard normal and then apply DWT), project the samples to the nearest CLWE point and then add noise in the secret direction. We can then reverse any transforms needed to get the samples in the required domain.
Input: Secret direction $\mathbf{w}$, Samples $\mathbf{y}_{1},\ldots,\mathbf{y}_{m}\in\mathbb{R}^{n}$
Output: Samples $\mathbf{y}^{\prime}_{1},\ldots,\mathbf{y}^{\prime}_{m}\in\mathbb{R}^{n}$
1 $\gamma^{\prime}\leftarrow\sqrt{\beta^{2}+\gamma^{2}}$;
2 for $i\in[m]$ do
3 $k_{i}\leftarrow\lceil\gamma^{\prime}\langle\mathbf{y}_{i},\mathbf{w}\rangle\rfloor$;
4 $z_{i}\leftarrow\mathcal{N}(0,\beta)$;
5 $\mathbf{y}_{i}^{\prime}\leftarrow\mathbf{y}_{i}+\left((z_{i}+k_{i}\cdot\gamma/\gamma^{\prime})/\gamma^{\prime}-\langle\mathbf{y}_{i},\mathbf{w}\rangle\right)\mathbf{w}$;
6
return $\mathbf{y}^{\prime}_{1},\ldots,\mathbf{y}^{\prime}_{m}$
Algorithm 1 Method to convert samples from a base distribution to hCLWE.
Algorithm 1 shows the details of this method, which is slightly more complicated due to the fact that the gaussian mixtures have width $\beta/\gamma^{\prime}=\beta/\sqrt{\beta^{2}+\gamma^{2}}$ and are separated by $\gamma/{\gamma^{\prime}}^{2}$. If $\beta\ll\gamma$ then $\gamma^{\prime}\approx\gamma$ and this simplifies to the expected steps. Indeed, for practical purposes the difference is insignificant. This can be seen in Figure 4, as the blue area is a histogram generated by a simulation of $10,000$ samples using this approximation and is not simply the area under the dashed line (the expected PDF).
Claim 3.1.
The method shown in Algorithm 1, on input from the standard normal distribution $\mathcal{N}$, produces samples from $H_{\mathbf{w},\beta,\gamma}$.
Proof.
Consider a single input sample $y$ drawn from the standard normal $\mathcal{N}$. Let $\gamma^{\prime}=\sqrt{\beta^{2}+\gamma^{2}}$, $\tilde{y}=\langle\mathbf{y},\mathbf{w}\rangle$ and $\mathbf{y}^{\perp}=\mathbf{y}-\tilde{y}\cdot\mathbf{w}$, i.e. $\tilde{y}$ is the projection of $\mathbf{y}$ in the secret direction and $\mathbf{y}^{\perp}$ is the component of $\mathbf{y}$ perpendicular to it. Set $k=\lceil\gamma^{\prime}\tilde{y}\rfloor$. We can now rewrite the output of the algorithm as:
$\mathbf{y}^{\prime}=\mathbf{y}^{\perp}+\left(\frac{z+k\cdot\gamma/\gamma^{\prime}}{\gamma^{\prime}}\right)\mathbf{w}=\mathbf{y}^{\perp}+\tilde{y}^{\prime}\mathbf{w}$
From this we can write the PDF of $\mathbf{y}^{\prime}$ as the convolution of the choices of $k$ and $z$:
$\rho(\mathbf{y}^{\perp})\cdot\sum_{k\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta}(z)$
By substituting out $z$, we get:
$\rho(\mathbf{y}^{\perp})\cdot\sum_{k\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta/\gamma^{\prime}}\left(\tilde{y}^{\prime}-k\gamma/{\gamma^{\prime}}^{2}\right)$
This expression can be manipulated algebraically to achieve the same expression as in Definition 1 (see Appendix B). ∎
5. Evaluation
Returning to the problem statement in Section 2.1 and specifically the model shown in Figure 2, we evaluate CLUE-Mark against the following questions:
  1. RQ1:
    What practical CLWE parameters should be used for CLUE-Mark?
  2. RQ2:
    How well can we recover the CLUE-Mark watermark after inverting the raw images?
  3. RQ3:
    What is the quality of the images generated by CLUE-Mark?
  4. RQ4:
    Do steganographic attacks work against CLUE-Mark?
  5. RQ5:
    How robust is CLUE-Mark to benign transformations, such as JPEG compression, on the watermarked images?
A proof-of-concept implementation of CLUE-Mark and source code to reproduce our evaluations is available at: https://github.com/kisp-nus/cluemark
5.1. RQ1. Practical Parameters and an Adaptive Covariance Attack
The CLWE problem is relatively recent candidate for a cryptographic hardness assumption, and as such, optimal parameters for security are not fully settled. The original reductions from hard problems to CLWE (Bruna et al., 2021) require that $\gamma\geq 2\sqrt{n}$ and a non-trivial value of $\beta>0$, using a quantum reduction. In (Gupte et al., 2022), it is further confirmed using a classical reduction that $\gamma=\tilde{\Omega}(\sqrt{n})$ is required for theoretical security. Using the stronger assumption of subexponential hardness of LWE, it is also shown in (Gupte et al., 2022) that $\gamma={\left(\log n\right)}^{\frac{1}{2}+\delta}\log\log n$ should be sufficient, for arbitarily small $\delta>0$. All of these are shown to be sufficient asymptotically. The question becomes: what concrete values of $n,\gamma$, and $\beta$ should we use in practice? In particular, image sampling and inversion processes add significant noise, so it is diserable to use the smallest possible $\gamma$ and $\beta$ to ensure that we can recover the watermark, while maintaining undetectability.
To answer this question, we extended the covariance attack described in (Bruna et al., 2021) to produce an adaptive attack that can work with fewer samples. In this attack, the adversary computes the covariance matrix of a large set of samples, computes the eigenvalues, and finds the maximum difference with $1/2\pi$, see Algorithm 2 for an explicit description. In the original attack, the adversary checks if this score is greater than $\gamma^{2}\exp(-\pi(\beta^{2}+\gamma^{2}))$. While this attack is successful when large number of samples is available, i.e. $m>O\left(\exp(\gamma^{2})\right)$, the threshold given is not useful for smaller sample sizes. For practical usage, we are interested in parameters that are secure for adversaries that have access to a more limited number of samples, i.e. $m$ at most polynomial in $\gamma$ or $n$. To that end, we adjust the procedure slightly to compare the maximum eigenvalues in the covariance matrix produced by a set of CLWE samples and normally distributed samples and use the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) as a measure of the attack’s success. This produces an attack that can be successful with fewer samples, at the expense of possibly more false positives.
Input: Samples $A\in\mathbb{R}^{m\times n}$
Output: Covariance score
1 $\Sigma_{m}\leftarrow\frac{1}{2\pi m}A^{T}A$;
2 $\mu_{1},\ldots,\mu_{n}\leftarrow\text{eigenvalues}(\Sigma_{m})$;
3 return $\max_{i\in[n]}\left|\mu_{i}-\frac{1}{2\pi}\right|$;
Algorithm 2 Covariance method to determine if samples are from a CLWE distribution or standard normal.
For each combination of $n$ (dimensionality), $m$ (number of samples), and $\gamma$ (gaussian pancake spacing), and fixed $\beta=0.001$, we produced $100$ iterations using both CLWE samples and normally distributed samples, and then score both using Algorithm 2. We then compute the AUC score, such that $1.0$ indicates that it is possible to distinguish the two sets, and $0.5$ indicates that the two sets are indistinguishable. A score in between indicates the presence of false positives or false negatives depending on the threshold used. We plot the results in Figure 8. As expected, with $\gamma=8$ (theoretically sound for $n\leq 64$) the attack is unsuccessful even with a million samples. At the opposite extreme, with $\gamma=1$ the attack is successful with several thousand samples. Interestingly, the attack requires more samples with larger $n$. This makes sense, as covariance estimation requires linearly more samples as the number of dimensions increases. If $\gamma$ is increased to $2$ or $4$, then in each case the number of samples needed for a successful attack increases an order of magnitude each time. Also plotted in Figure 8 are the attack accuracies given by the threshold recommended by prior theoretical work (Bruna et al., 2021). As expected, for smaller $\gamma$ values the theoretical and practical attacks converge quickly. For larger $\gamma$ values, the thresholds become less useful, and were unable to produce any distinguishing power for any $\gamma$ greater than $1$.
In the CLWE hardness literature (Bruna et al., 2021; Gupte et al., 2022), the security parameter is usually defined as the number of dimensions ($n$), with security proven for $\gamma\geq\Omega(\sqrt{n})$, implying that $\gamma$ should grow with $n$ to maintain security when the number of samples $m$ is not fixed. However, when the sample size is fixed, our results show that increasing $n$ while keeping $\gamma$ fixed may actually improve security. In particular, if the total amount of data ($n\cdot m$) is fixed, then increasing $n$ while keeping $\gamma$ fixed reduces $m$, which makes the problem harder.
[…]
Figure 8. Covariance attack scores from statistical simulations. Each combination of $n,m$ and $\gamma$ was simulated $100$ times for both CLWE and Normal distributions to produce scores using Algorithm 2. The resulting ROC AUC is shown on the vertical axis. A score of 0.5 is equivalent to random guessing, and 1.0 indicates that it’s possible to perfectly distinguish the two categories. The theoretical attack accuracy (”t. acc.” entries) correspond to the accuracy of using the threshold value $\gamma^{2}\exp(-\pi(\beta^{2}+\gamma^{2}))$ as suggested in the original work (Bruna et al., 2021).
In the context of diffusion model images, the latent space is typically $64\times 64\times 4=2^{14}$. Thus, if we divide the latent space into blocks of size at least $n\geq 32$, this produces $m\leq 512$ samples per image, well under 1,000 samples. Based on the results of our test above, we conclude that $\gamma$ as small as $2$ should be sufficient.
With $n\geq 32$, setting $\gamma\geq 2$ and $\beta\geq 0.001$ should provide practical indistinguishability of CLWE from the normal distribution, provided the number of samples is limited to $m\leq 1,000$.
We stress that the above parameters are based on our empirical evaluations and not on the theoretical reductions of CLWE hardness given in prior works. Further research on tighter reductions and the concrete parameters to use in practice is important future work. For this reason we provide theoretical proofs of security that apply to any choice of $\gamma$ and $\beta$, but use the parameters from this section in order to show a proof-of-concept.
5.2. Evaluation Setup
We perform our evaluation on state-of-the-art text to image open sourced diffusion models, specifically, Stable Diffusion version 2.1. We generate $512\times 512$ images starting with latent embeddings of size $4\times 64\times 64$ in $50$ steps. We use the inversion process of (Hong et al., 2024) for obtaining latent vector estimates from iamges.
Datasets.
We use two standard benchmark datasets: Stable Diffusion Prompts (SDP) [^5] https://huggingface.co/datasets/Gustavosta/Stable-Diffusion-Prompts, and COCO (Lin et al., 2014). We generate an image without a watermark and for each watermarking technique using the same prompt and initial seed, for each of the first $100$ prompts in each dataset.
Baselines.
Tree Ring and Gaussian Shading are the two state-of-the-art pre-processing based watermarking techniques that claim some level of undetectability unlike the post-processing based methods which are known to be detectable (Wen et al., 2023). We use these two methods as our baselines. We use the “ring” pattern for Tree Ring, and the standard Gaussian Shading method, but note that we fixed the key and nonce for all images. The original code for Gaussian Shading generated a fresh key and nonce for each image, which is not practical since the recovery process would not have access to the key and nonce for a specific image. All parameters for Tree Ring and Gaussian Shading match the defaults published in the respective papers.
Evaluation Metrics.
We use Frechet Inception Distance (FID) (Heusel et al., 2017) to measure the quality of generated images. FID measures the difference in the embeddings between two images using the Inception model, which is a heuristic for how similar they are. Measuring the FID between the watermarked images and their unwatermarked counterparts gives a sense of the quality degradation of the watermarking scheme, i.e. a lower FID implies better quality images.
All of the watermarking schemes produces a score indicative of the likelihood of the watermark being present in the image. Thus, the performance of the watermarking scheme depends on choosing a threshold for this score. This produces a true positive rate (TPR, corresponds to completeness in Definition 3) and false positive rate (FPR, corresponds to soundness in Definition 4) for a given threshold. By plotting the TPR vs FPR curve for varying thresholds we obtain the Receiver Operating Characteristic (ROC) curve, a commonly used method for evaluating the performance of binary classifiers. The area under the ROC curve (AUC) is a metric that summarizes the performance of the classifier. A score of 1 indicates perfect performance, while a score of 0.5 indicates performance no better than random. Thus we use the AUC score to evaluate the performance of the watermarking schemes.
System Specification.
We ran our experiments on an AMD EPYC 7443P 4GHz processor with 96 GB of RAM and 4 Nvidia A40 GPUs with 45 GB of memory each, running CUDA 12.2 and Nvidia drivers 535.129.03.
5.3. RQ2. Watermark Recovery
We answer how sound and complete CLUE-Mark’s recovery strategy is by producing a watermarked and unmarked image for each of the $100$ prompts. As discussed in Section 5.1 we set $\gamma=2,\beta=0.001$ with a block size of $2\times 4\times 4$ an apply the watermark in the DWT domain. We then apply CLUE-Mark’s recovery strategy to all the images and compare the results. We find that for both datasets, CLUE-Mark is able to distinguish between watermarked and unmarked images with high accuracy, with an AUC score greater than 0.99 in both cases. For comparison, the baseline techniques produced AUC scores of $1.0$ for the same conditions.
CLUE-Mark’s detector has high soundness and completeness (AUC scores $\geq 0.99$) in unperturbed images.
5.4. RQ3. Image Quality
As can be seen in Figure 9, CLUE-Mark generates images with the closest quality to the non-watermarked images for both datasets, while Gaussian Shading and Tree Ring produces an FID distance significantly greater (worse) than CLUE-Mark.
The degradation in quality of the images indicates that both the baselines are not undetectable, especially when many watermarked images are analyzed at once for detectability. This is expected for Tree Ring since it does not embed provably undetectable watermarks. It is also expected for Gaussian Shading as it only proves undetectability for a single image. We explain this in detail in our discussion regarding Gaussian Shading in Section 5.5.
CLUE-Mark generates images with similar quality to non-watermarked images and significantly higher quality images than the baselines.
COCOSDP05050100100150150144.04144.04142.8142.8180.99180.99172.68172.6840.340.342.1442.14FID ScoreTree RingGaussian ShadingCLUE-Mark
Figure 9. Image quality as measured by FID to unwatermarked images on two standard datasets. Lower numbers indicate higher quality (closer to unwatermarked).
5.5. RQ4. Resilience to Steganographic Attacks
A steganographic attack attempts to remove the watermark from marked images by averaging the difference between the watermarked and unwatermarked images over many samples. In order to compare with prior work and confirm the theory behind our construction, we reproduced the results from (Yang et al., 2024a) using $100$ pairs of watermarked and unwatermarked images generated for each watermarking method, and average the difference between the images. As can be seen in the examples in Figure 11, there are obvious patterns in the averages from Tree Ring and Gaussian Shading, but CLUE-Mark produces a nearly blank image. We then subtract this average from the watermarked images and attempt to recover the watermark. As shown in Figure 10, the attack successfully removes the Tree Ring and Gaussian Shading watermarks, but is unable to remove the CLUE-Mark watermark. We note that the resulting watermark-removed images for Tree Ring are of similar quality, but for Gaussian Shading suffer some degradation. We expect that a more sophisticated attack could remove the Gaussian Shading watermark without degrading the image.
CLUE-Mark is robust to standard steganographic attacks unlike the baselines whose watermark can be detected and removed.
COCOSDP0.40.40.60.60.80.8110.510.510.430.430.510.510.580.580.980.980.980.98AUCTree RingGaussian ShadingCLUE-Mark
Figure 10. AUC scores after steganographic attack. Values closer to 1 indicate presence of the watermark after the attack, while closer to 0.5 indicate removal of the watermark.
Figure 11. Steganographic attack examples against related schemes. All images generated using the prompt “a dog laying next to some colorful blankets on a white bed” with the same random seed. Gaussian shading is evaluated with a fixed nonce.
5.6. RQ5. Robustness to Benign Transformations
We evaluate the robustness of CLUE-Mark and the baseline techniques under different non-adversarial post-processing techniques. Specifically, we apply JPEG compression (quality value $0.95$), brightness (level 1), and random cropping (0.9 in each direction), and rotations to the watermarked images and measure recovery performance. Our results are shown in Figure 12. For comparison, the baselines all produced AUC scores at or near 1.0 for all conditions. We observe that CLUE-Mark is somewhat robust to JPEG compression and brightness changes, but not to the other perturbations. Thus CLUE-Mark is robust to minor perturbation, but any significant modifications make the watermark unrecoverable.
CLUE-Mark is robust to minor perturbations such as JPEG compression and brightness but not robust to larger perturbations.
JPEGCropBrightness0.40.40.60.60.80.8110.970.970.540.540.70.70.960.960.50.50.750.75AUC ScoreCOCOSDP
Figure 12. Robustness to common perturbations as measured by RoC AUC for CLUE-Mark. Values closer to 1.0 indicate watermark presence after perturbations, while closer to 0.5 indicates removal.
6. Related Work
Watermarking images has been extensively studied since the 1990s (Van Schyndel et al., 1994). Early research primarily focused on watermarking individual images, while more recent advancements target AI-generated images. Unlike traditional images, AI-generated content allows watermarking at various stages of the generation pipeline, including altering training data or fine-tuning models to embed watermarks.
Individual Images.
Post-processing techniques are commonly applied to individual images, embedding watermarks by modifying images after creation. Classical methods operate in either the spatial (Van Schyndel et al., 1994; Tsai et al., 2000) or frequency domain (Al-Haj, 2007; Guo and Georganas, 2002; Hamidi et al., 2018; Kundur and Hatzinakos, 1997; Lee et al., 2007). Spatial domain techniques modify pixel intensities, often by altering the least significant bits of selected pixels. While efficient, these techniques lack robustness against simple transformations like brightness adjustments or JPEG compression. In contrast, frequency domain methods embed watermarks by altering an image’s frequency coefficients after applying a transformation such as the Discrete Cosine Transform (DCT) or Discrete Wavelet Transform (DWT). These approaches offer greater robustness to perturbations, while ensuring that the watermark remains imperceptible, often measured by peak signal-to-noise ratio (PSNR).
More recently, many deep learning-based techniques have been introduced in order to enhance robustness and quality (Jia et al., 2021; Tancik et al., 2020; Zhu, 2018; Liu et al., 2019; Ahmadi et al., 2020). These techniques typically leverage encoder-decoder neural architectures, where the encoder embeds a message into an image, and the decoder recovers the message. The models are trained to minimize differences between the watermarked and original images for even an adversarial classifier, while ensuring message recovery even in the presence of noise. Various architectures, noise injection strategies, and training objectives have been explored to achieve robustness against diverse transformations.
However, existing post-processing methods lack provable guarantees of undetectability and may degrade image quality (Wen et al., 2023; Fernandez et al., 2023). Additionally, using the same watermark across multiple images can expose vulnerabilities to steganographic attacks, allowing the watermark to be extracted (Yang et al., 2024a).
Generative Models.
Generative models like Generative Adversarial Networks (GANs) and Latent Diffusion Models have become standard for producing high-fidelity, ultra-realistic images. When traditional watermarking methods are applied to AI-generated images, they often reduce visual quality (Fernandez et al., 2023), which is undesirable.
Recent research addresses this issue by embedding watermarks directly within the generative process (Yu et al., 2022, 2021; Fernandez et al., 2023; Wen et al., 2023; Yang et al., 2024b; Lukas and Kerschbaum, 2023). One approach involves training parts of the model to embed messages during generation, allowing extraction via a pre-trained decoder (Yu et al., 2022; Fernandez et al., 2023; Yu et al., 2021; Lukas and Kerschbaum, 2023). However, these techniques require multiple training rounds, which is computationally expensive, especially as the size of state-of-the-art models continues to grow. While these methods claim improved image quality over post-processing techniques, they rely on ad-hoc metrics like PSNR and are significantly more resource-intensive.
Watermarking Diffusion Models.
Recent advances in watermarking latent diffusion model images propose modifying only the input latent vectors, offering computational efficiency without altering model parameters (Wen et al., 2023; Yang et al., 2024b). Diffusion models then generate images normally, unaware of the modification to the latents. The watermarks are detected by reversing the diffusion process and examining latent vector. Techniques such as Tree Ring (Wen et al., 2023) and Gaussian Shading (Yang et al., 2024b) claim state-of-the-art performance in both visual quality, in the sense that they are imperceptible to human observers, and robustness, in the sense that after common image perturbations the watermark can still be recovered. However, none of these techniques provide provable undetectability; Tree Ring provided no proof of undetectability only that it was claimed to be “imperceptible”, while the proof of Gaussian Shading undetectability requires a per-image nonce that must be available at verification time, an assumption that is difficult to realize in practice. Moreover, both watermarks can still be identified, removed, or forged by steganographic attacks (Yang et al., 2024a).
In work concurrent to ours, (Gunn et al., 2025) proposes using Pseudo-Random Codes (PRC) to bias the noise input to a diffusion model, similar to Gaussian Shading but without requiring a nonce. As the name suggests, PRCs are a new cryptographic primitive that can be considered a combination of encryption and error-correcting codes. Cyphertexts are generated by encoding a message with a secret key, which can then be decoded even if some subset of the bits are flipped. Without the secret key the cyphertext appears random to all computationally-bounded observers. Intuitively, the signs of the latent noise are chosen to match a PRC encoding of a 0-length message. By the pseudorandom nature of the PRC, this is equivalent to choosing the signs randomly, and therefore does not affect the model output. Similar to Tree Ring and Gaussian Shading, the model is inverted to obtain an estimate of the original noise vector, from which the sign values are extracted and checked using the PRC decoder. If the decoder outputs a valid message, the image is considered watermarked. In a sense, one could see CLUE-Mark as a continuous-domain PRC using CLWE, in that with the key the presence of the code can still be detected despite noise, while without the key the samples are indistinguishable from a Gaussian. Interestingly, in the appendix of (Gunn et al., 2025), the authors describe that in order to achieve robustness to perturbations they had to use parameters that are not theoretically secure. The parameters picked in CLUE-Mark also do not match those of in prior theoretical reductions, but for a completely different reason: We are building on the security of CLWE against a sample-constrained adversary, which has not been theoretically studied thus far to our knowledge.
Overall, our work advances these approaches by introducing a watermarking scheme that is based on CLWE, which is an alternative cryptographic construct. We aim for computational undetectability is our main goal, while addressing robustness against non-adversarial transformation (such as image compression) that is common when transmitting images over the web or social networks. Our tool, CLUE-Mark, achieves good image quality in latent diffusion model outputs yet watermarks cannot be identified or forged without the secret key in our empirical evaluation. Ours is a novel application of CLWE in the context of generative models, as prior research using CLWE focused on implanting undetectable backdoors in random Fourier feature-based classification models (Goldwasser et al., 2022).
7. Conclusions and Future Work
In this work we introduced CLUE-Mark, a provably undetectable watermarking scheme for diffusion model generated images, leveraging the cryptographic hardness of the Continuous Learning With Errors (CLWE) problem. We first defined formally what it means for a watermark scheme to be undetectable, and then described how CLUE-Mark works. We then proved that CLUE-Mark is undetectable under the hCLWE assumption with limited samples. Finally, we evaluated empirically the hCLWE parameter choices against common attacks, and the performance of CLUE-Mark on common diffusion models and datasets, showing that it is usable in practice, produces images with much higher fidelity to the original unwatermarked images than existing techniques, and cannot be detected by steganographic techniques. CLUE-Mark is robust to some common perturbations, such as JPEG compression and brightness changes. The parameters we chose provide a practical security level.
While CLUE-Mark presents the first demonstration of how CLWE-based watermarks are already usable for diffusion models, it can be improved. Future work can consider improving the robustness of CLUE-Mark to perturbations. The concrete parameters for CLWE are not yet well studied, and it is possible that better parameters could improve the performance of CLUE-Mark while still maintaining undetectability. Finally, the technique used for inverting the model to obtain the estimate of the original latent vector is the current bottleneck in recovering the watermark in terms of both quality and speed. Improvements to the inversion method could also significantly improve the performance of CLUE-Mark.
Appendix A Are Perfect Watermarks possible?
Ideally, a watermark should be robust in the sense that perturbations that do not affect the image quality should not remove the watermark. Unfortunately, it is impossible for a watermark to simultaneously satisfy completeness, soundness, provable undetectability, and perfect robustness. We formalize the statement and provide a proof.
Undetectability.
Recall that in Definition 6 we defined undetectability as the property that no PPT algorithm can distinguish between a watermarked image and its unwatermarked counterpart, without access to the watermarking key $k$. In other words, the quality of the distribution of the images before and after watermarking does not change, from the perspective of any quality oracle that does not have access to keys and is computationally bounded. Lemma 1 bounds the distance between distributions observed under two keys $k,k^{\prime}$ against any such quality oracle. Informally, this tells us that the quality of images generated from watermarking, even with different keys, is the same to any downstream function without the watermarking keys.
Lemma 1.0.
For an undetectabile watermarking scheme, no PPT distinguisher $D$ can distinguish between $\mathcal{W}_{\theta,k}$ and $\mathcal{W}_{\theta,k^{\prime}}$ with non-negligible advantage. That is, $Adv_{D,\mathcal{W}_{\theta,k},\mathcal{W}_{\theta,k}}\leq negl(\lambda)$, assuming $D$ has no access to the keys $k$ or $k^{\prime}$.
Proof.
The proof is by a standard hybrid argument. Consider any PPT distinguisher $D$. From the definition of undetectability, we have that $Adv_{D,\mathcal{M}_{\theta},\mathcal{W}_{\theta,k}}\leq negl(\lambda)$ and $Adv_{D,\mathcal{M}_{\theta},\mathcal{W}_{\theta,k^{\prime}}}\leq negl(\lambda)$. By the triangle inequality:
$Adv_{D,\mathcal{W}_{\theta,k},\mathcal{W}_{\theta,k^{\prime}}}$ $\leq Adv_{D,\mathcal{M}_{\theta},\mathcal{W}_{\theta,k}}+Adv_{D,\mathcal{M}_{\theta},\mathcal{W}_{\theta,k^{\prime}}}$
$\leq 2\cdot negl(\lambda)=negl(\lambda).$
Perfect Robustness.
If we allow arbitrary perturbations of the image, it is trivial to remove the watermark by simply replacing the image with another completely different image. What may not be immediately apparent is that even if we restrict to perturbations that maintain image quality, it is still impossible to have a watermark that is robust to such perturbations and is undetectible. Here we define quality in the same way as undetectability: the quality of two sets of images is the same if no PPT algorithm can distinguish the distributions with non-negligible advantage.
A watermarking scheme is perfectly robust if a watermarked image, after perturbations that do not change its quality, is treated the same as the unperturbed image by the (keyed) verifier, i.e., the $Extract$ function accepts the perturbed image if the watermarked image before perturbation was accepted. More formally, a perfectly robust watermark is one wherein given an $x\leftarrow\text{Mark}(k,\pi)$, no PPT adversary can find an image $x^{\prime}$ of the same quality as $x$ but for which $\text{Extract}(k,x)\neq\text{Extract}(k,x^{\prime})$.
Theorem 2.
No watermarking scheme satisfying soundness, completeness, undetectability, and perfect robustness exists.
Proof.
Assume for the sake of contradiction that a watermarking scheme satisfying all four properties exists. We can instantiate the scheme with a key $k\leftarrow Setup(1^{\lambda})$ and by computing $x\leftarrow Mark(k,\pi)$ for an input $\pi$. By the completeness property, the $Extract(k,x)$ must accept. Now consider the following generic attack strategy: The adversary who does not know $k$ generates a fresh key $k^{\prime}\leftarrow Setup(1^{\lambda})$ and generates $x^{\prime}\leftarrow Mark(k^{\prime},\pi)$. Note that by Lemma 1, PPT algorithms without access to $k$ and $k^{\prime}$ cannot distinguish the distributions of $x$ and $x^{\prime}$ with non-negligible probability. Thus, $x$ and $x^{\prime}$ have the same quality. We analyze the behavior of the original verifier, i.e., the distribution of the random variable $Extract(k,x^{\prime})$:
  • If the verifier $Extract(k,x^{\prime})$ accepts with probability more than $negl(\lambda)$, then it violates soundness. This is because $k^{\prime}\neq k$ with high probability $1-negl(\lambda)$, and with that probability, we have a PPT adversary who does not know $k$ but has found $x^{\prime}$ that is accepted by the verifier with the key $k$.
  • If the verifier $Extract(k,x^{\prime})$ rejects with probability more than $negl(\lambda)$, then the scheme violates perfect robustness—the adversary has efficiently  [^6] The adversary breaks the hardness criterion while sticking to the constraint stated for perfect robustness, i.e., knowing $k^{\prime}$ but not $k$. found $x^{\prime}$ of same quality  [^7] Note that quality is determined by oracles with no access to keys. as $x$ (by Lemma 1) without knowing $k$, and yet we have $Extract(k,x^{\prime})\neq Extract(k,x)$.
We have derived a contradiction to the assumptions in both cases above, which are exhaustive and mutually exclusive. Thus, no watermarking scheme satisfies all 4 properties. ∎
The attack strategy underlying our proof of Theorem 2 is realizable within practical threat models, such as those assumed in recent attacks on LLM watermarking schemes (Pang et al., 2024).
Appendix B HCLWE Algebra
First we notice the following equivalence:
$\left(\frac{k}{\gamma^{\prime}}\right)^{2}+\left(\frac{\tilde{y}^{\prime}-k\gamma/{\gamma^{\prime}}^{2}}{\beta/\gamma^{\prime}}\right)^{2}$
$=\frac{k^{2}}{\beta^{2}+\gamma^{2}}+\frac{\beta^{2}+\gamma^{2}}{\beta^{2}}\left(\left(\tilde{y}^{\prime}\right)^{2}-\frac{2\gamma\tilde{y}^{\prime}k}{\beta^{2}+\gamma^{2}}+k^{2}\frac{\gamma^{2}}{\left(\beta^{2}+\gamma^{2}\right)^{2}}\right)$
$=\left(\tilde{y}^{\prime}\right)^{2}+\frac{\left(\gamma\tilde{y}^{\prime}\right)^{2}}{\beta^{2}}-\frac{2\left(\gamma\tilde{y}^{\prime}\right)k}{\beta^{2}}+\frac{k^{2}}{\beta^{2}}$
$=\left(\tilde{y}^{\prime}\right)^{2}+\left(\frac{\gamma\tilde{y}^{\prime}-k}{\beta}\right)^{2}$
Where in the second step the last term comes from noticing that:
$\frac{k^{2}}{\beta^{2}+\gamma^{2}}+\frac{\gamma^{2}}{\beta^{2}}\frac{k^{2}}{\beta^{2}+\gamma^{2}}=\frac{k^{2}}{\beta^{2}+\gamma^{2}}\left(1+\frac{\gamma^{2}}{\beta^{2}}\right)=\frac{k^{2}}{\beta^{2}}$
We can now rearrange the statement in the claim as:
$\rho(\mathbf{y}^{\perp})\cdot\sum_{k\in\mathbb{Z}}\rho_{\gamma^{\prime}}(k)\rho_{\beta/\gamma^{\prime}}\left(\tilde{y}^{\prime}-k\gamma/{\gamma^{\prime}}^{2}\right)$
$=\exp\left(-\pi\|\mathbf{y}^{\perp}\|^{2}\right)\cdot\sum_{k\in\mathbb{Z}}\exp\left(-\pi\left(\frac{k}{\gamma^{\prime}}\right)^{2}\right)\exp\left(-\pi\left(\frac{\tilde{y}^{\prime}-k\frac{\gamma}{{\gamma^{\prime}}^{2}}}{\beta/\gamma^{\prime}}\right)^{2}\right)$
$=\sum_{k\in\mathbb{Z}}\exp\left[-\pi\left(\|\mathbf{y}^{\perp}\|^{2}+\left(\frac{k}{\gamma^{\prime}}\right)^{2}+\left(\frac{\tilde{y}^{\prime}-k\frac{\gamma}{{\gamma^{\prime}}^{2}}}{\beta/\gamma^{\prime}}\right)^{2}\right)\right]$
$=\sum_{k\in\mathbb{Z}}\exp\left[-\pi\left(\|\mathbf{y}^{\perp}\|^{2}+\left(\tilde{y}^{\prime}\right)^{2}+\left(\frac{\gamma\tilde{y}^{\prime}-k}{\beta}\right)^{2}\right)\right]$
$=\rho(\mathbf{y}^{\prime})\sum_{k\in\mathbb{Z}}\rho_{\beta}(k-\gamma\langle\mathbf{w},\mathbf{y}^{\prime}\rangle)$
Where in the last step we used the orthogonality of $\mathbf{y}^{\perp}$ and $\mathbf{w}$ to write $\|\mathbf{y}^{\prime}\|^{2}=\|\mathbf{y}^{\perp}\|^{2}+\left(\tilde{y}^{\prime}\right)^{2}$, and the definition of $\tilde{y}^{\prime}$.